OpenStudy (anonymous):

HELP please! What are the solutions of 3x2 + 4x = -5?

OpenStudy (anonymous):

Are we trying to write as standard form 3x^2+4x+5=0?

OpenStudy (anonymous):

the quantity negative 4 plus or minus i square root of 66 all over 6 the quantity negative 2 plus or minus 2i square root of 66 all over 3 the quantity negative 2 plus or minus i square root of 11 all over 3 the quantity negative 4 plus or minus i square root of 11 all over 6

OpenStudy (anonymous):

these are the answer choices

OpenStudy (anonymous):

@WAkawaka12

OpenStudy (zpupster):

\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array}

OpenStudy (zpupster):

you can try putting the values into the quadratic equation, above a=3,b=4,c=5

OpenStudy (zpupster):

\begin{align} \\ \frac{-b \pm \sqrt{ b^2 -4( a)(c)} }{2(a)}\\ \frac{-4\pm \sqrt{4^2 -4(3)(5)} }{2(3)}\\ \frac{-4\pm \sqrt{-44 } }{6} = \frac{-4\pm \sqrt{-4\times11 } }{6} \\ = \frac{-4\pm 2i\sqrt{11 } }{6} \\ = \frac{-2\pm 1i\sqrt{11 } }{3} \ \end{align}

OpenStudy (zpupster):

here is an example --practice with a few of these --lot of work but not so bad.